1. A nitrogen molecule can be thought of as two point masses (m of each = 14u, where u = 1.67 x 10-27 kg) separated by a distance of 1.3 x 10-10 m. In air at room temperature the average rotational kinetic energy of such a molecule is about 4.0 x 10-21 J. Find the moment of inertia of such a molecule about its center of mass, and its speed of rotation in revolutions per second. 
1
Expert's answer
2022-02-23T12:56:54-0500

"I=\frac{md^2}2=19.76\cdot 10^{-47}~kg\cdot m^2,"

"E=\frac{I\omega^2}2,"

"\omega=\sqrt{\frac{2E}I},"

"\nu=\frac{\omega}{2\pi}=\sqrt{\frac E{2\pi^2 I}}=3.2\cdot 10^{17}~\frac{rev}s."